# Are Open-Vocabulary Models Ready for Detection of MEP Elements on Construction Sites

### v1

@@ 1 Introduction
Figure 1: Our Journeybot platform for data collection: a Jackal rover base from Clearpath equipped with several cameras and LiDARs.
@@
This paper evaluates the potential of open vocabulary object detection models for MEP monitoring in construction environments. To this end, we constructed a real-world dataset of MEP components using a mobile robot equipped with multiple sensors. The dataset was used to train and evaluate a fine-tuned supervised object detection model, YOLO11 [12], and compare its performance with three open vocabulary object detection models: Grounding-SAM2 [13], Grounding DINO [14], and DETIC [15]. This comparative analysis provides a comprehensive assessment of fine-tuned, lightweight closed-set detectors versus pre-trained open-vocabulary models in real-world MEP monitoring tasks. Our findings offer valuable insights into the effectiveness of deploying open-vocabulary models compared to smaller, task-specific models for dynamic construction environments. As LLMs and open-vocabulary models continue to evolve, this work contributes to understanding their role in advancing construction site automation.
@@ 2 Experimental Setup
The sensing system was tailored for comprehensive environmental monitoring and digital twin data collection. It included four Intel Realsense D435i stereo cameras covering all sides of the robot, with one facing upward, and a front-facing Intel Realsense T265 tracking camera for enhanced localization. Depth images from the D435i supported collision avoidance, together with a horizontally mounted Velodyne Puck 32MR LiDAR. An Ouster LiDAR augmented the data collection system, but both LiDAR are not covered by this work. Wheel encoder data and IMU readings were synchronized with detection sensors by an NVidia Jetson computer.
Dataset: The dataset was collected at an active construction site provided by an industrial partner, with a primary focus on MEP assets. Using the Robot Operating System (ROS), our robotic platform was manually teleoperated through the site, capturing data stored as rosbags. Images were extracted at a rate of 10 Hz from the rosbags, followed by a filtering process to remove redundant and blurred frames. The remaining images were annotated for the 10 specific MEP components listed in Table 1 using an online annotation tool.
Table 1: Dataset details with number of images (and instances) per each dataset split per each class.
@@ 10 Pump 1536(1844)
The final annotated dataset was divided into 70% training, 20% validation, and 10% testing splits. Table 1 also presents the number of images and instances per class. Example images are shown in the results section (Figure 3), illustrating the challenges posed by real-world data collection on a mobile platform. These include objects appearing at varying distances, being partially occluded, or blurred due to motion, reflecting the dynamic and unpredictable nature of construction site environment
Evaluation Procedure: For quantitative evaluation, we assessed the performance of a lightweight object detector (YOLO11 Nano pre-trained on COCO [16]) fine-tuned (200 epochs) on our dataset, comparing its best performing weights against three pre-trained large open-vocabulary models without fine-tuning: Grounding DINO, Grounded SAM2, and DETIC. The evaluation metrics included precision, recall, and F1 score (All measured at 0.5 intersection over union) which served as qualitative measures of performance.
@@ 4 Conclusion
Future research could focus on narrowing this gap by fine-tuning open-vocabulary models on specialized datasets, potentially combining their adaptability with the precision of supervised models, to better address the challenges of dynamic construction environments.

### v2

@@ 1 Introduction
Figure 1: Our Journeybot platform for data collection: a Jackal rover base from Clearpath equipped with several cameras and LiDARs (not used in this study).
@@
This paper evaluates the potential of open vocabulary object detection models for MEP monitoring in construction environments. To this end, we constructed a real-world dataset of MEP components using a mobile robot equipped with multiple sensors. The dataset was used to train and evaluate a fine-tuned supervised object detection model, YOLO11 [12], which as a closed-set detector works only for the specific classes that it was trained or fine-tuned on. Then, compare its performance with three open vocabulary object detection models: Grounding-SAM2 [13], Grounding DINO [14], and DETIC [15]. This comparative analysis provides a comprehensive assessment of fine-tuned, lightweight closed-set detectors versus pre-trained open-vocabulary models in real-world MEP monitoring tasks. Our findings offer valuable insights into the effectiveness of deploying open-vocabulary models compared to smaller, task-specific models for dynamic construction environments. As LLMs and open-vocabulary models continue to evolve, this work contributes to understanding their role in advancing construction site automation.
@@ 2 Experimental Setup
The sensing system was tailored for comprehensive environmental monitoring and digital twin data collection. It included four Intel Realsense D435i stereo cameras covering all sides of the robot, with one facing upward, and a front-facing Intel Realsense T265 tracking camera for enhanced localization. Depth images from the D435i supported collision avoidance, together with a horizontally mounted Velodyne Puck 32MR LiDAR. An Ouster LiDAR augmented the data collection system, but both LiDARs are not covered by this work. Wheel encoder data and IMU readings were synchronized with detection sensors by an NVidia Jetson computer.
Dataset: The dataset was collected at an active construction site provided by an industrial partner, with a primary focus on MEP assets. Using the Robot Operating System (ROS), our robotic platform was manually teleoperated through the site, capturing data stored as rosbags. Images were extracted at a rate of 10 Hz from the rosbags, followed by a filtering process to remove redundant and blurred frames. The remaining images were annotated for the 10 specific MEP components listed in Table 1 using an online annotation tool. The dataset is shared online.
Table 1: Dataset details with number of images (and instances) per each class.
@@ 10 Pump 1536(1844)
The final annotated dataset was divided into 70% training, 20% validation, and 10% testing splits. Table 1 also presents the number of images and instances per class. Example images are shown in the results section (Figure 3), illustrating the challenges posed by real-world data collection on a mobile platform. These include objects appearing at varying distances, being partially occluded, or blurred due to motion, reflecting the dynamic and unpredictable nature of construction site environments.
Evaluation Procedure: For quantitative evaluation, we assessed the performance of a lightweight object detector (YOLO11 Nano pre-trained on COCO [16]) fine-tuned (200 epochs) on our dataset, comparing its best performing weights (on the validation split) against three pre-trained large open-vocabulary models without fine-tuning: Grounding DINO, Grounded SAM2, and DETIC. The evaluation metrics included precision, recall, and F1 score (All measured on the testing split at 0.5 intersection over union).
@@ 4 Conclusion
With the huge advancements in vision-language models, we hope this work encourages research on narrowing this gap, possibly, by fine-tuning open-vocabulary models on specialized datasets, potentially combining their adaptability with the precision of supervised models. Another direction could be to design construction-related prompts to adapt the open-vocabulary detectors.